OpenStudy (anonymous):

What is the solution set of (x - 2)(x - 3) = 2? a) {1, 4} b) {2, 3} c) {4, 5}

OpenStudy (anonymous):

(x - 2)(x - 3) = 2 x^2-3x-2x+6=2 x^2-5x+6=2 x^2-5x+6-2=0 x^2-5x-4=0 dELTA=25-16=9 x1=5-3=1 x2=5+3=4 S={1,4} hope it's correct :)